media%2Ff25%2Ff25a5a55-1caf-432f-9144-c0
Answer #1

9 balls at 12.30

9+9 balls at 12.45

9*3 balls at 12.525

9*4 balls at 12.5625

9*5 balls at 12.58125

9*6 balls at 12.590625

9*7 balls at 12.5953

9*8 balls at 15.59.76

and soon

After 12.59 the interval becomes less than one minute which will get further divided by 2. This number is less than the interval between 12.59 and 1 This happens the next time for eg Initially you have 30 minutes to remove the ball it can be removed before 12.30. The interval between 12.30 and 1.00 is 30 minutes (Say t).In the next experiment the time to remove reduces to 15 minutes (lets say this as s).This 15 minutes will be added to 12.30 . This happens every time ie a half of the time left is getting added every time ie t>s always. That is why the process will never go beyond the time 1,00

9(1+2+3....+\infty)=\infty For both of them there are infinitely many balls in their barrel

The answers do not make sense. as it quite impossible to drop and remove the balls in micro or nano seconds.